write an equation for the ellipse passing through (0, 2) and (4, 0) with center at the origin
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)
Answer by stanbon(75887)

write an equation for the ellipse passing through (0, 2) and (4, 0) with center at the origin
--------------------
Form: x^2/a^2 + y^2/b^2 = 1
------
(0,2) gives: 0 + 4/b^2 = 1
b^2 = 4
-----
(4,0) gives: 16/a^2 + 0 = 1
a^2 = 16
-------------
Equation:
x^2/16 + y^2/4 = 1
========================
Cheers,
Stan H.